OpenStudy (anonymous):

Can some one plz help me in this question find the exact gradient of the tangent to the curve y=e^x+ln x at the point where x=1 I know y'= 1/x times e^x+lnx And then sub 1 But i am not getting answer of 2e

OpenStudy (anonymous):

times?

OpenStudy (anonymous):

As in * multiply

OpenStudy (anonymous):

\[y=e^x+\ln(x)\] \[y'=e^x+\frac{1}{x}\]

OpenStudy (anonymous):

X+lnx are both power

OpenStudy (anonymous):

e^(x+lnx)

OpenStudy (anonymous):

oooh it is \[e^{x+\ln(x)}=e^xe^{\ln(x)}=xe^x\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

Srry i am on my ipad

OpenStudy (dan815):

back to PEDMAS with you

OpenStudy (anonymous):

your derivative is wrong

OpenStudy (anonymous):

Oh hw so

OpenStudy (dan815):

yah its 1+1/x dummy!

OpenStudy (anonymous):

first off \[\large e^{x+\ln(x)}=e^xe^{\ln(x)}=xe^x\] and to take the derivative of that, use the product rule there is no \(\frac{1}{x}\) in it

OpenStudy (anonymous):

e^x = e^x and e^lnx will be y'= 1/x e^lnx

OpenStudy (dan815):

^ or you can keep it as the exponential and do d/dx ( e^(fx))= f'x * e^f(x) d/dx (e^(x+lnx))=(1+1/x)*e^(x+lnx)

OpenStudy (dan815):

what ever you like best

OpenStudy (anonymous):

The derivative of e^x is not x

OpenStudy (anonymous):

whoa hold the phone \[\frac{1}{x}e^{\ln(x)}=1\]

OpenStudy (anonymous):

Hahah thank u

OpenStudy (dan815):

you are welcome ! :)

OpenStudy (anonymous):

there is really no reason ever to write \(e^{\ln(x)}\) since it is just \(x\)

OpenStudy (anonymous):

I know but In hsc our exam they cut off marks for that